Aqueous hydrobromic acid HBr

 will react with solid sodium hydroxide NaOH

 to produce aqueous sodium bromide NaBr

 and liquid water H2O

. Suppose 61. g of hydrobromic acid is mixed with 38.4 g of sodium hydroxide. Calculate the maximum mass of sodium bromide that could be produced by the chemical reaction. Be sure your answer has the correct number of significant digits.


1
Expert's answer
2021-09-20T02:26:32-0400

HBr + NaOH = NaBr + H2O

According to the equation, n (NaBr) = n (HBr) = n (NaOH).

n = m / M

M (NaBr) = 102.9 g/mol

M (HBr) = 80.9 g/mol

M (NaOH) = 40 g/mol

n (HBr) = 61 / 80.9 = 0.75 mol

n (NaOH) = 38.4 / 40 = 0.96 mol

Therefore, HBr is the limiting reactant.

n (NaBr) = n (HBr) = 0.75 mol

m (NaBr) = n x M = 0.75 x 102.9 = 77.2 g